Solve the following equation:
[0 0 -1]
[3 x 0 ] = x^2
[2 0 3]

Answer by Alan3354(69443)

[0 0 -1]
[3 x 0 ] = x^2
[2 0 3]
------------
0 - 0 -1(0-2x) = x^2
2x = x^2
x = 0
x = 2